# zkSTAR: A zero knowledge system for time series attack detection enforcing regulatory compliance in critical infrastructure networks

### v1

@@ Abstract
Industrial control systems (ICS) form the operational backbone of critical infrastructure networks (CIN) such as power grids, water supply systems, and gas pipelines. As cyber threats to these systems escalate, regulatory agencies are imposing stricter compliance requirements to ensure system-wide security and reliability. A central challenge, however, is enabling regulators to verify the effectiveness of detection mechanisms without requiring utilities to disclose sensitive operational data. In this paper, we introduce zkSTAR, a cyberattack detection framework that leverages zk-SNARKs to reconcile these requirements and enable provable detection guarantees while preserving data confidentiality. Our approach builds on established residual-based statistical hypothesis testing methods applied to state-space detection models. Specifically, we design a two-pronged zk-SNARK architecture that enforces (i) temporal consistency of the state-space dynamics and (ii) statistical consistency of the detection tests—allowing regulators to temporally verify alarm correctness without visibility into utility-level data. We formally analyze the soundness and zero-knowledge properties of our framework and validate its practical feasibility through computational experiments on real-world ICS datasets. As a result, our work demonstrates a scalable, privacy-preserving alternative for regulatory compliance for ICS driven critical infrastructure networks.
I Introduction
@@
Critical Infrastructure Networks (CINs)—such as gas, water, and power systems—are large-scale cyber-physical systems managed by multiple stakeholder entities. The growing integration of information and operational technologies (IT-OT) driven by increased IoT adoption has introduced new vulnerabilities to data-driven cyberattacks. Such attacks often target the Industrial Control Systems (ICS) of utility stakeholders, aiming to degrade asset performance and potentially trigger cascading failures across the entire network. Although attack detection is performed locally at stakeholder ICSs, regulatory entities such as Information Sharing and Analysis Centers (ISACs) aggregate alarms and statistics from all utilities in real time to enhance situational awareness, reduce false positives, and expedite decision-making. The effectiveness of ISACs in fulfilling their objectives fundamentally depends on the quality of local ICS detection metrics reported by individual utilities. Inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions. Consequently, there remains a critical need for methods that allow regulators to publicly verify local ICS detection outcomes without resorting to exhaustive utility data audits, which are costly, impractical for real-time operations, and intrusive to utility privacy. Therefore, this paper introduces zkSTAR, a framework that enables publicly verifiable detection of data-driven ICS attacks while preserving complete zero knowledge of local ICS state-space dynamics.
The ICS serves as the primary OT backbone for any CIN utility capable of providing control, actuation and data storage capabilities of infrastructure assets in a scalable manner. Data-driven ICS attack detection methods primarily utilize models that characterize the state space of the local utility network which can help unveil deviations from expected steady state conditions using the observed sensor measurements [1]. These deviations or residuals can then be used as inputs for statistical hypothesis tests to distinguish between normal and abnormal operation profiles subject to Type-I and Type-II error parameters of the test [2]. An alarm signifying an attack event is triggered locally when the test reject its null hypothesis indicating statistical inconsistencies between the observed and the expected sensor measurements subject to state space dynamics.
With existing state of the art methods, the only way to conduct regulatory verification of detection outcomes of utility-based state space models is through a full-fledged data audit of ICS operations. For such an audit, the utility would have to relinquish the ownership of their local IoT datasets to a regulator raising significant privacy and data governance hurdles. Next, utility datasets must be shown to be temporally consistent across the time domain. In other words, the regulator must be able to exhaustively certify the state space estimates of successive time steps have been generated from previous estimates. Finally, the regulator must also be able to demonstrate statistical consistency which implies that the correct statistical testing procedure was followed to raise alarms. As a result, regulatory data audits can become extremely expensive and would come at the cost of significant privacy issues for utilities.
@@
Therefore, in this paper, we develop zkSTAR, a framework that is geared towards guaranteeing temporal and statistical consistency for state space driven statistical detection of data driven ICS attacks. At its core, zkSTAR is specifically designed to demonstrate temporal and statistical consistency of ICS systems with highly non-linear state space dynamics in complete zero knowledge. zkSTAR is based on the zkSNARK (zero-knowledge succinct, non-interactive argument of knowledge) paradigm and provides the ability to publicly verify attack detection outcomes for a state space modeling paradigm in a trustworthy fashion. We present a detailed overview of our contributions as follows:
@@
  • We establish comprehensive theoretical guarantees pertaining to the necessary and sufficient conditions for asserting temporally and statistically consistent regulatory compliance in the context of time-series anomaly detection using non-linear state space models.
@@
  • We design a zkSNARK based kernel decomposition scheme that provides the ability to prove temporal and statistical consistency claims in a computationally efficient manner.
@@
  • We provide a detailed container-based computational testbed for evaluating the performance of zkSTAR with respect to a detailed case studies involving real-world datasets.
II Related Works
Attacks on ICSs that are solely targeted towards the IT systems can be detected and isolated using network traffic monitoring paradigms [3, 4, 5]. However, data-driven attacks are harder to detect through such mechanisms especially in the ICS context owing to their ability to impact critical pathways in both the IT and OT layers [6] potentially causing critical infrastructure damage. Such attacks typically involve manipulation of sensor data so as to effect cascading failures across networked ICS in addition to degrading the performance of critical infrastructure assets. Despite robust network traffic monitoring frameworks, such attacks can go undetected since they involve manipulation at the OT layer[7, 8].
Therefore, several model-based mechanisms have been proposed to detect ICS data-driven attacks [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Among these, the Kalman filter based mechanism of detecting attack proves to be a robust and flexible option [2, 1, 22, 23, 24] owing to its reliance on residuals emanating from a well-formed state space model. Such techniques utilize the residuals obtained from Kalman model predictions and the observed sensor measurements as fundamental inputs to statistical hypothesis tests in order to flag alarms. As a result, a Kalman filter based state space estimation mechanism can help detect anomalies emanating from data driven attacks distinguish between routine faults and attacks using degradation models [22] and use decentralized methodologies to raise network-wide alarms [23].
From the regulatory perspective however, it is essential to not only receive timely notifications regarding attack detection at utility stakeholders but also ensure the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates. Given the trepidation of utilities in being able to share data with regulatory agencies due to privacy and logistical hurdles, enabling regulatory compliance of local detection frameworks acquires an entirely new degree of complexity. As a result, it becomes important to ensure that the right sequence of steps has been followed to determine these insights without belying the underlying data. Zero-knowledge proofs form a viable alternative to solve this conundrum with the added capability of complementing privacy mechanisms as well [25].
Zero-knowledge proofs were originally proposed by Goldwasser et al. [26] in which a Prover attempts to convince a Verifier that some mathematical statement is true. The basic premise was that the Verifier must not glean any other information about the secrets held by the Prover that validate a set of mathematical statements. Original propositions of zero-knowledge proofs were interactive, i.e. the Verifier and the Prover exchange a series of messages in order to establish the validity of the mathematical truths. However, Blum et al. [27] showed that zero-knowledge proofs can also be made non-interactive by using a common short random string (CRS). These initial works were followed by establishing the concepts of succinctness that enables efficient verification with minimal overhead [28] which were combined with non-interactivity [29].
In the past decade, there have been several breakthroughs that have paved the way for practical, computationally feasible implementations of zero-knowledge Succinct, Non-Interactive Arguments of Knowledge (zk-SNARKs) [30, 31, 32, 33, 34]. Within the domain of machine learning, zero-knowledge proofs have been used to establish the integrity of training [35, 36, 37] and prove model-ownership [38]. However, there exists a critical gap pertaining to the development of zkSNARKs for time-series anomaly detection using state space dynamics of utility ICS. As a result, considering state of the art methods, providing trustworthy data driven attack detection for fulfilling regulatory compliance requirements in complete zero-knowledge in currently infeasible. In this paper, we develop zkSTAR, a framework that can generate computationally efficient zkSNARKs of data driven attack detection frameworks that are based on Kalman filter driven utility ICS state space models. zkSTAR is primarily geared towards ensuring regulatory compliance in data driven attack detection with full privacy preservation of utility stakeholder data. In fact, zkSTAR is capable of enabling regulatory agencies to verify utility level attack detection mechanisms in zero knowledge of their local state space model altogether.
III Nonlinear State Space Modeling
@@
For developing zkSTAR, we leverage the extended Kalman Filter state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [39]. Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge. Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
III-A Extended Kalman Filter Model
@@
We present a brief overview of the Extended Kalman Filter model in this section while a detailed description of the model has been provided in Appendix -A. Let g,h{g},{h} denote the state transition and observation functions respectively, $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors. Additionally, let process and measurement noises follow multivariate normal distributions $N(0,Q_{t})$ and $N(0,R_{t})$ respectively. Now, we define our EKF model based on the following equations:
@@
III-B Hypothesis Test for Anomaly Detection
We consider $r_{t}$ as our residual and $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ as the residual covariance. We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [1, 40]. This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [40].
@@ III-B Hypothesis Test for Anomaly Detection
A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in Appendix -B.
We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@
IV Regulatory Compliance using zkSNARKs
Detecting attacks requires monitoring covariance matrices of the residuals to drive the underlying $\chi^{2}$ hypothesis tests. However, publicly validating alarms by an external entity such as a regulator would require the disclosure of residual vectors and covariance matrices that enable these hypothesis tests. Therefore, a critical regulatory challenge is the verification of the hypothesis test outcomes as reported by the utilities in the absence of their state space residual and covariance estimates. Zero-knowledge proofs form a viable alternative for solving these critical pain points by helping achieve public verification of utility level attack detection outcomes for regulatory purposes.
Therefore, in this paper, we employ zero knowledge proofs to enable regulatory compliance and verification. We concentrate on a utility-oriented subsystem represented by a state-space model as outlined in Section III. Our privacy considerations specifically address the disclosure risks linked to the public revelation of parameters $C$ and residuals $r_{t}$ derived from the Kalman Filter-based state-space estimation, which are crucial for validating the alarm value $\rho_{t}$. As a result, our methodology specifically leverages zk-SNARKs to help assert temporal consistency and statistical consistency for utility-level $\chi^{2}$ hypothesis tests without the need to divulge state space variables.
IV-A Zero Knowledge Preliminaries
@@
Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates. We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [41].
@@
The zk-SNARK proof generation phase can be defined by function Prove such that
@@
Definition 5 details the proving phase of the zk-SNARK. The Prove function consumes the proving key $pk$, $\mathcal{M}$ parametrized by $\theta$ as well as the inputs $\Delta^{in}_{t}$ and outputs $\Delta^{out}_{t}$ at time $t$ to generate a proof artifact $\Pi_{t}$. The Prove function needs to be executed for each timestep $t$ at the utility level.
@@
Zero knowledge regulatory compliance for attack detection relies on two fundamental operational attributes of each utility. These fundamental operational attributes pertain to temporal consistency and statistical consistency. We argue that these two fundamental operational attributes are necessary and sufficient conditions to enable zero-knowledge regulatory compliance verification. Our argument is based on the ability of the pretrained model $\mathcal{M}$ to recursively ensure that state space estimates are derived from previous predictions while being statistically consistent with expected residual covariance. In the following subsections, we formally state each operational attribute separately as a precursor for guaranteeing integrity of detection outcomes.
IV-B Temporal Consistency
Temporal consistency pertains to the recursive relationship between a posteriori state space estimates at $t$ and $t+1$ respectively. Temporal consistency constraint encoded in the non-linear Kalman filter model ensures that the posteriori estimate $\hat{x}_{t+1|t+1}$ is a derivative of $\hat{x}_{t|t}$. We state this formally as part of Definition 7.

### v2

@@ Abstract
Industrial control systems (ICS) form the operational backbone of critical infrastructure networks (CIN) such as power grids, water supply systems, and gas pipelines. As cyber threats to these systems escalate, regulatory agencies are imposing stricter compliance requirements to ensure system-wide security and reliability. A central challenge, however, is enabling regulators to verify the effectiveness of detection mechanisms without requiring utilities to disclose sensitive operational data. In this paper, we introduce zkSTAR, a cyberattack detection framework that leverages zk-SNARKs to reconcile these requirements and enable provable detection guarantees while preserving data confidentiality. Our approach builds on established residual-based statistical hypothesis testing methods applied to state-space detection models. Specifically, we design a two-pronged zk-SNARK architecture that enforces (i) temporal consistency of the state-space dynamics and (ii) statistical consistency of the detection tests, enabling regulators to verify correctness and prevent suppression of alarms without visibility into utility-level data. We formally analyze the soundness and zero-knowledge properties of our framework and validate its practical feasibility through computational experiments on real-world ICS datasets. As a result, our work demonstrates a scalable, privacy-preserving alternative for regulatory compliance for ICS driven critical infrastructure networks.
1 Introduction
@@
The convergence of information and operational technologies (IT–OT) through widespread IoT adoption has expanded the attack surface for data-driven cyber intrusions. Such attacks often target industrial control systems (ICSs), degrading asset performance and triggering cascading failures across critical infrastructure networks (CINs). While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [dhs_oig2]. The effectiveness of these regulators, however, hinges on the integrity and reliability of local detection data. Recent real-world incidents [langner2011stuxnet, jeffries2022cyber, salazar2024tale] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies. Such incidents expose a fundamental regulatory blind spot pertaining to verifying the integrity of reported detections without re-auditing raw operational data. Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [dhs_oig2]. To address this gap, we introduce zkSTAR, a framework enabling publicly verifiable detection of data-driven ICS attacks in complete zero knowledge of local ICS state-space dynamics.
zkSTAR addresses a critical regulatory gap by enabling provable temporal and statistical consistency verification of ICS detections. Leveraging knowledge-soundness and provably consistent outcomes, it ensures that compromised utilities cannot falsify or suppress alarms even if their IT–OT stacks are breached. The framework enables verifiable regulatory oversight while preserving utility data confidentiality [dhs_oig2], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [nolan2015cybersecurity]. Additionally, by adopting an on-demand, lazy proof generation paradigm, zkSTAR ensures a low utility overhead as well.
Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [li2020detection]. Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [li2020deep]. With existing approaches, regulators can verify such detection outcomes only through full data audits of ICS operations—requiring utilities to relinquish proprietary IoT datasets and raising severe privacy and data-governance concerns. Moreover, utilities must demonstrate both temporal consistency wherein successive state estimates are derived from prior ones; and statistical consistency in which alarms follow valid hypothesis test procedures. These audits are costly, slow, and incompatible with real-time oversight.
@@
To overcome these limitations, zkSTAR delivers provable temporal and statistical consistency guarantees for state space driven statistical detection of data driven ICS attacks. At its core, zkSTAR is specifically designed to validate consistency claims for ICS systems with highly non-linear state space dynamics in complete zero knowledge. zkSTAR is based on the zkSNARK (zero-knowledge succinct, non-interactive argument of knowledge) paradigm providing publicly verifiable attack detection outcomes for a state space modeling paradigm in a trustworthy fashion. zkSTAR operates in an on-demand compliance mode, where regulators issue proof-generation requests to utilities whenever verification of local detections is required. We present a detailed overview of our contributions as follows:
@@
  • We establish formal security guarantees for zkSNARK driven regulatory compliance with temporal and statistical consistency proofs of detection outcomes that also ensure resistance to attack-suppression under knowledge soundness property.
@@
  • We design a zkSNARK based kernel decomposition scheme that provides the ability to prove temporal and statistical consistency claims in a scalable and computationally efficient manner.
@@
  • We detail the performance of zkSTAR with respect to detailed case studies involving real-world datasets using a REST-API driven container-based computational testbed.
Overall, we show that zkSTAR enforces integrity of detection proofs under adversarial falsification, preventing utilities from reusing stale states or forging alarms—making it a security enforcement layer for regulatory trust.
2 Related Works
Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [caselli2016specification, ye2004robustness]. However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [urbina2016limiting, kang2014cyber, drias2015taxonomy], potentially causing cascading failures across critical infrastructure. More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [salazar2024tale], Triton [jeffries2022cyber] and Stuxnet [langner2011stuxnet]. These vulnerabilities have demonstrated the feasibility of joint manipulation of control systems and monitoring data which have motivated the need for robust verifiable detection mechanisms for ICS.
To address some of these challenges, several model-based mechanisms have been proposed [huang2018online, hoehn2016detection, rahman2017multi]. Among these, Kalman filter based approaches are particularly robust and flexible [li2020deep, li2020detection, li2021degradation, ramanan2021blockchain, li2022online], leveraging residuals from state-space models as statistical evidence for anomaly detection. These residual-based tests differentiate routine faults from malicious attacks [li2021degradation] and can support decentralized, network-wide alarm dissemination [ramanan2021blockchain].
From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [ramanan2021blockchain, dhs_oig2]. Therefore, regulatory compliance goals would ideally enable stakeholders to provide publicly verifiable detection outcomes that are accurate and trustworthy without exposing private operational data. Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [narayan2015verifiable], enabling a prover to convince a verifier of correctness without revealing sensitive information.
Originating from foundational work by Goldwasser et al. [goldwasser2019knowledge] and Blum et al. [blum2019non], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [kilian1992note, micali2000computationally, gennaro2013quadratic, parno2016pinocchio, ben2013snarks, costello2015geppetto, ben2014succinct]. Recent applications in machine learning verify training integrity and model ownership [huang2022zkmlaas, wang2022ezdps, nguyen2022preserving, yang2023fedzkp]. While existing zk-SNARK frameworks enable general-purpose verifiable computation, none target the verification of dynamical system integrity or hypothesis-test correctness which are cornerstones of industrial attack detection. Recent works on verifiable differential privacy [narayan2015verifiable] and secure multiparty analytics [froelicher2020drynx] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses. To alleviate this core critical gap, this paper introduces zkSTAR, a framework that enables provable, privacy-preserving verification of Kalman filter–based detection mechanisms, paving the way for regulatory compliance in CINs.
3 Nonlinear State Space Modeling
@@
For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [coskun2017long]. Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge. Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
3.1 Extended Kalman Filter Model
@@
We present a brief overview of the EKF model in this section while a detailed description of the model has been provided in Appendix A.1. Let g,h{g},{h} denote the state transition and observation functions respectively, $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors. Additionally, let process and measurement noises follow multivariate normal distributions $N(0,Q_{t})$ and $N(0,R_{t})$ respectively. Now, we define our EKF model based on the following equations:
@@
3.2 Hypothesis Test for Anomaly Detection
We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [li2020detection]. Without loss of generality we note that it is also possible to apply our techniques for detecting false data injection by monitoring $R_{t}$. We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [li2020detection, hwang2017chi]. This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [hwang2017chi].
@@ 3.2 Hypothesis Test for Anomaly Detection
A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [li2021degradation]. We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@
4 Regulatory Compliance using zkSNARKs
Detecting attacks requires monitoring covariance matrices of the residuals to drive the underlying $\chi^{2}$ hypothesis tests. However, publicly validating alarms by an external entity such as a regulator would require the disclosure of residual vectors and covariance matrices that enable these hypothesis tests. Therefore, a critical regulatory challenge is the verification of the hypothesis test outcomes as reported by the utilities in the absence of their state space residual and covariance estimates. Zero-knowledge proofs form a viable alternative for solving these critical pain points by helping achieve public verification of utility level attack detection outcomes for regulatory purposes. Therefore, in this paper, we employ zero knowledge proofs to enable regulatory compliance and verification. We concentrate on a utility-oriented subsystem represented by a state-space model as outlined in Section 3 that is meant to be captured in a zkSNARK framework. As a result, our methodology specifically leverages zk-SNARKs to assert temporal consistency and statistical consistency for utility-level $\chi^{2}$ hypothesis tests in zero-knowledge of local state space dynamics.
4.1 Zero Knowledge Preliminaries
@@
Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates. We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [menezes1996handbook].
@@
The zk-SNARK proving phase is defined by function Prove such that
@@
Definition 5 details the proving phase of the zk-SNARK. The Prove function consumes the proving key $pk$, $\mathcal{M}$ parametrized by $\theta$ as well as the inputs $\Delta^{in}_{t}$ and outputs $\Delta^{out}_{t}$ at time $t$ to generate a proof artifact $\Pi_{t}$. The Prove function is executed for each timestep $t$ at the utility level.
@@
Zero knowledge regulatory compliance for attack detection relies on two fundamental operational attributes of temporal and statistical consistency. We argue that these two fundamental operational attributes are necessary and sufficient conditions to enable zero-knowledge regulatory compliance verification. Our argument is based on the ability of the pretrained model $\mathcal{M}$ to recursively ensure that state space estimates are derived from previous predictions while being statistically consistent with expected residual covariance. In the following subsections, we formally state each operational attribute separately as a precursor for guaranteeing integrity of detection outcomes.
4.2 Temporal Consistency
Temporal consistency (TC) pertains to the recursive relationship between a posteriori state space estimates at $t$ and $t+1$ respectively. Temporal consistency constraint encoded in the non-linear Kalman filter model ensures that the posteriori estimate $\hat{x}_{t+1|t+1}$ is a derivative of $\hat{x}_{t|t}$. We state this formally as part of Definition 7.